OpenStudy (anonymous):

On planet X, an object weighs 18.2 N. On planet B where the magnitude of the free-fall acceleration is 2.39 g (where g = 9.8 m/s^2 is the gravitational acceleration on Earth), the object weighs 32.7 N. The acceleration of gravity is 9.8 m/s^2. What is the mass of the object be on Earth?

OpenStudy (shamim):

On planet B Weight w=32.7N a=2.39g=2.39*9.8=? m=?

OpenStudy (shamim):

Use the formula w=ma 32.7=m*2.39*9.8 m=?

OpenStudy (shamim):

Response plz!!!!!!

OpenStudy (anonymous):

m=32.7/(2.39*9.8)=1.396 g

OpenStudy (shamim):

Not g but kg. Right??

OpenStudy (anonymous):

yes

OpenStudy (shamim):

U should give me medal?????!!!!! @imyint

OpenStudy (anonymous):

sorry i didnt realize that was the answer

OpenStudy (anonymous):

how do i find the free fall acceleration of planet x tho?